# Towards Efficient Training of Graph Neural Networks: A Multiscale Approach

### v3

@@ 4.3 Multiscale Gradients Computation
In Section 4.1Section 4.2 we presented two possible methods to facilitate the efficient training of GNNs via a multiscale approach. In what follows, we present a complementary method that builds on these ideas, as well as draws inspiration from Monte-Carlo optimization techniques (Giles, 2001), to further reduce the cost of GNN training.
We start with the observation that optimization in the context of deep learning is dominated by stochastic gradient-based methods (Kingma & Ba, 2014; Ruder, 2016). At the base of such methods, stands the gradient descent method, where at optimization step $s$ we have:
@@
for some $01>\gamma_{r}>\gamma_{r-1}>\ldots>\gamma_{1}=$ 0. To understand why this concept is beneficial, let us consider the discretization of the fine scale problem (i.e., for $\mathcal{G}_{1}$) in Equation (4) using $M$ samples. Suppose that the error in approximating the expected loss at the first, finest scale ${\mathbb{E}}[\mathcal{L}_{s}^{(1)}]$ by the samples behaves as $e^{(1)}={\frac{C}{\sqrt{M}}}$. Now, let us consider a 2-scale process, with $M_{1}$ samples on the finest resolution graph. The error in evaluating the loss on the second scale ${\mathbb{E}}[\mathcal{L}_{s}^{(2)}]$ with $M_{2}$ examples is given by $e^{(2)}={\frac{C}{\sqrt{M_{2}}}}$ while the error in evaluating ${\mathbb{E}}[(\mathcal{L}_{s}^{(1)}-\mathcal{L}_{s}^{(2)})]$ is $e^{(1,2)}={\frac{C\gamma_{1}}{\sqrt{M_{1}}}}$. Since the assumption is that $\gamma<1$, we are able to choose $M_{2}$ and $M_{1}$ such that $e^{(1)}\approx e^{(2)}+e^{(1,2)}$. Hence, rather than using $M$ finest graph resolution observations, we use only $M_{1}<M$ such observations and sample the coarser problems instead.
@@
${\mathbb{E}}\,\mathcal{L}^{(1)}_{s}({\boldsymbol{\theta}})$$\approx$ $\textstyle{{\frac{1}{M_{R}}}\sum\,\mathcal{L}^{(R)}_{j}({\boldsymbol{\theta}})}+\textstyle{{\frac{1}{M_{R-1}}}\sum\,\left(\mathcal{L}^{(R-1)}_{s}({\boldsymbol{\theta}})-\mathcal{L}^{(R)}_{s}({\boldsymbol{\theta}})\right)}+...$
$+\textstyle{{\frac{1}{M_{1}}}\sum\,\left(\mathcal{L}^{(1)}_{s}({\boldsymbol{\theta}})-\mathcal{L}^{(2)}_{s}({\boldsymbol{\theta}})\right)},$
where $M_{R}>M_{R-1}>M_{1}$. This implies that for the computation of the loss on the fine resolution $\mathcal{G}_{1}=\mathcal{G}$, only a few fine graph approximations are needed, and most of the computations are done on coarse, small graphs. Our method is summarized in Algorithm 2, and we provide an illustration of the computation process in Figure 3. Naturally, this method suits inductive learning tasks, where multiple graphs are available.
@@
Figure 3: Illustration of the Multiscale Gradient Computation algorithm introduced in Section 4.3.
@@ 4.4 A Motivating Example
Table 1: Experiments for different coarsening strategies, creating coarse graph with $\frac{1}{4}$ of the original nodes. Loss discrepancy between the loss computed on the original graph, and the pooled graph ($\Delta$loss) is calculated using $\gamma_{s}$ in Equation (9), and $p$ denotes the power in Equation (5).
Coarsening type ##\## edges $\Delta$loss (initial) $\Delta$loss (final)
@@
To explore this further, we train a four-layer GCN Kipf & Welling (2016) on a dataset of 100 rods in varying orientations and evaluate different coarsening methods for constructing a reduced graph. Table 1 presents the loss values for various coarsening strategies along with the corresponding edge counts. The results highlight another aspect of our approach: methods that do not increase connectivity prior to pooling (i.e., with $p=1$) may, in certain situations, lead to significant variance in loss across different coarsened graphs. For this dataset, introducing additional connectivity before pooling successfully mitigates this issue while still reducing the graph size in terms of both nodes and edges.
To further evaluate the capabilities of our multiscale methods on this synthetic dataset, Appendix B presents the full multiscale training results and compares them with standard fine-grid training. Additionally, we construct a second synthetic dataset derived from the MNIST image dataset LeCun et al. (1998). Results on both datasets demonstrate that our methods effectively capture global graph features while enabling efficient training.
@@ 4.5 Theoretical Properties of Reduced Graphs
At the base of all the above discussed techniques, the core is the ability of the coarse graph to represent the fine graph. To this end, consider the model problem
${\ell}({\boldsymbol{\Theta}})={\frac{1}{2N}}\|\sigma({\bf A}{\bf X}{\boldsymbol{\Theta}})-{\bf Y}\|^{2},$ (11)
and assume that $\sigma$ is the identity mapping. One way to analyze the solution of the coarse-scale problem is through sketching (see Woodruff (2014)). In sketching, the original problem is replaced by
${\ell}_{c}({\boldsymbol{\Theta}})={\frac{1}{2N}}\|{\bf P}^{\top}{\bf A}{\bf X}{\boldsymbol{\Theta}}-{\bf P}^{\top}{\bf Y}\|^{2}$ (12)
where ${\bf P}^{\top}$ is a sketching matrix. We establish the following result:
@@
Let ${\boldsymbol{\Theta}}^{*}$ be the solution of equation (11) and let ${\boldsymbol{\Theta}}_{C}^{*}$ be the solution of equation (12). Let ${\bf P}^{\top}$ be a sparse subspace embedding with $O(\frac{c^{2}}{\epsilon})$ rows, where $c$ is the size of the feature vectors (number of graph channels), that injects any vector into a subset of indices, which we denote by $C$. According to this subset, denote ${\bf A}=\begin{bmatrix}{\bf A}_{CC}&{\bf A}_{CF}\\ {\bf A}_{FC}&{\bf A}_{FF}\end{bmatrix}$, and ${\bf X}=\begin{bmatrix}{\bf X}_{C}\\ {\bf X}_{F}\end{bmatrix}$. Then if $\sigma=Id$, there exists $\|{\bf R}\|<\|{\bf A}_{CF}{\bf X}_{F}{\boldsymbol{\Theta}}_{C}^{*}\|$ such that with probability .99:
${\ell}_{c}({\boldsymbol{\Theta}}_{C}^{*})={\frac{1}{2N}}\|{\bf P}^{\top}{\bf A}{\bf P}{\bf P}^{\top}{\bf X}{\boldsymbol{\Theta}}_{C}^{*}-{\bf P}^{\top}{\bf Y}\|^{2}\leq{\frac{1\pm\epsilon}{2N}}\left(\|{\bf A}{\bf X}{\boldsymbol{\Theta}}^{*}-{\bf Y}\|^{2}+\|{\bf R}\|^{2}\right)$ (13)
The proof is provided in Appendix A.
@@ 5 Computational Complexity
As noted in Section 3, the computational cost of GNN message-passing layers is directly influenced by the size of both the node feature matrix and the adjacency matrix. This observation suggests that reducing the number of nodes and edges can lead to improved computational efficiency during training.
To illustrate this, consider Graph Convolutional Networks (GCNs) Kipf & Welling (2016), where each layer follows the update rule defined in Equation 1. When analyzing floating-point operations (FLOPs), we account for both the input and output dimensions of each layer, as well as the sparsity of the adjacency matrix, meaning that only nonzero entries contribute to computation. The total FLOPs required for a single GCN layer is given by:
@@
where $\cal{E}$ represents the set of edges, $\cal{V}$ the set of nodes, $c_{in}$ the input feature dimension, and $c_{out}$ the output feature dimension of the layer. From this expression, it follows that reducing the number of nodes and edges leads to a decrease in computational cost. For instance, if the number of nodes is halved, i.e., $|{\cal{V_{C}}}|=\frac{1}{2}|{\cal{V}}|$, and the number of edges satisfies $|{\cal{E_{C}}}|<|{\cal{E}}|$ for any pooling strategy, then the FLOPs per layer are reduced accordingly. This reduction translates into a significant theoretical improvement in efficiency.
Importantly, this effect is not limited to GCNs. A similar computational advantage is observed for other GNN architectures, like Graph Isomorphism Networks (GIN) Xu et al. (2018) and Graph Attention Networks (GAT) Velickovic et al. (2017). In all cases, reducing the graph size results in fewer computations per layer, demonstrating a generalizable advantage across different GNN architectures.
@@ 6 Experiments
We demonstrate that our approach achieves comparable or superior downstream performance to standard GNN training on the original graph while reducing computational costs, across the training techniques outlined in Section 4 and the pooling strategies discussed below. We demonstrate our methods on both mid-size and large-scale datasets, showing that given a need to solve a certain graph problem, our methods will prove to be efficient given existing computational resources.
@@
Our method incorporates various coarsening techniques to enhance GNN training efficiency:
@@
    Topk Coarsening: Selects $n_{r}$ nodes with the highest degrees before computing ${\bf A}_{r}$ using Equation (5). This approach may be computationally expensive for large graphs as it requires full degree information.
@@
    Subgraph Coarsening: Samples a random node $j$ and selects its nearest $n_{r}$ neighbors. If geometric positions are available, the selection is based on Euclidean distance; otherwise, a $k-hop$ ego network is used. This method captures local graph structures but may overlook global properties.
@@
    Hybrid Coarsening: Alternates between Random and Subgraph Coarsening to leverage both high- and low-frequency components of the graph.
@@ 6.1 Transductive Learning
We evaluate our approach using the OGBN-Arxiv dataset Hu et al. (2020) with GCN Kipf & Welling (2016), GIN Xu et al. (2018), and GAT Velickovic et al. (2017). This dataset consists of 169,343 nodes and 1,166,243 edges. We report test accuracy across three runs with different random seeds and include standard deviation to assess robustness, and present our results in Table 2, demonstrating that our multiscale framework achieves comparable or superior performance relative to standard training. Notably, random pooling proves to be the most effective approach for this dataset, offering both efficiency and strong performance. Given the large number of edges in this dataset, we use $p=1$ and report the corresponding edge counts in Table 3. The results highlight the substantial reduction in edges at each pooling level, significantly improving training efficiency. In Table 2, we also provide a timing analysis, including both theoretical FLOP estimates (as elaborated in Section 5) and empirical runtime measurements, confirming that our multiscale approach enhances efficiency while maintaining strong predictive performance. To show the efficiency of our methods, we demonstrate a comparison of the convergence of the training loss between the multiscale training and fine grid training in Appendix C.
@@
We demonstrate our results using the OGBN-MAG dataset Hu et al. (2020), which is a heterogeneous network composed of a subset of the Microsoft Academic Graph (MAG). This dataset includes papers, authors, institutions, and fields of study, of which the papers are labeled. We focus on the ’paper’ nodes and the ’citing’ relation, using a homogeneous graph dataset of 736,389 nodes and 5,416,271 edges. We summarize our results in Table 4, where we show that using GCN Kipf & Welling (2016), our methods achieve results comparable to the baseline results. We demonstrate our results for 2, 3, and 4 levels of multiscale training, and note that even though GCN is widely used and may be applied to this dataset, other more advanced architectures may yield better results than the presented baseline. However, changing a chosen model will not harm the effectiveness of our proposed methods.
@@
To emphasize that our method is general and will prove to be effective over various datasets and methods, we use additional datasets of various sizes and show that for every one of them, our methods prove to be efficient while yielding comparable or even better results compared to the original baseline. In Appendix D, we show results and details for the well-known Cora, SiteCeer, and PubMed datasets Sen et al. (2008). We further test our methods on Flickr Zeng et al. (2019), WikiCS Mernyei & Cangea (2020), DBLP Bojchevski & Günnemanng (2017), the transductive versions of Facebook, BlogCatalog, and PPI Yang et al. (2020) datasets, detailed in Appendix E. We present results for Flickr and PPI, along with timing evaluations, as well as full results for all those datasets and additional experimental details.
@@ 6.2 Inductive Learning
We evaluate our method on the ShapeNet dataset Chang et al. (2015), a point cloud dataset with multiple categories, for a node segmentation task. Specifically, we present results on the ’Airplane’ category, which contains 2,690 point clouds, each consisting of 2,048 points. Since this dataset contains multiple graphs, its overall size is large, and achieving an improvement in training efficiency for each batch will yield a significant impact on training efficiency.
In this dataset, nodes are classified into four categories: fuselage, wings, tail, and engine. For training, we use Dynamic Graph Convolutional Neural Network (DGCNN) Wang et al. (2018), which dynamically constructs graphs at each layer using the $k$-nearest neighbors algorithm, with $k$ as a tunable hyperparameter. We test our multiscale approach with random pooling, subgraph pooling, and a combination of both. In subgraph pooling, given that the data resides in $\mathbb{R}^{3}$, we construct subgraphs by selecting a random node and its $n_{j}$ nearest neighbors. Experiments are conducted using three random seeds, and we report the mean Intersection-over-Union (IoU) scores along with standard deviations in the top section of Table 5. Our results show that for all tested values of $k$, applying either random or subgraph pooling improves performance compared to standard training while reducing computational cost. Furthermore, combining both pooling techniques consistently yields the best results across connectivity settings, suggesting that this approach captures both high- and low-frequency components of the graph while maintaining training efficiency. Additional results for multiscale training on the full ShapeNet dataset are provided in Appendix F.
We apply the Multiscale Gradients Computation method to the ShapeNet ’Airplane’ category, evaluating various pooling techniques using Algorithm 2 with two levels. As shown in the bottom section of Table 5, our approach reduces training costs while achieving performance comparable to fine-grid training. These results show that the early stages of optimization can be efficiently performed using an approximate loss, as we propose, with fine-grid calculations needed only once the loss function nears its minimum. Additional results, timing analyses and experimental details for the ShapeNet dataset are provided in Appendix F.
@@
We further evaluate our Multiscale Gradients Computation method on the protein-protein interaction (PPI) dataset Zitnik & Leskovec (2017), consisting of graphs representing different human tissues. The results, with additional details, are provided in Appendix G.

### v5

@@ 4.3 Multiscale Gradients Computation
In Section˜4.1Section˜4.2, we presented two possible methods to facilitate the efficient training of GNNs via a multiscale approach. In what follows, we present a complementary method that builds on these ideas, as well as draws inspiration from Monte-Carlo optimization techniques (mlcm), to further reduce the cost of GNN training.
We start with the observation that optimization in the context of deep learning is dominated by stochastic gradient-based methods (kingma2014adam; ruder2016overview). At the base of such methods, stands the gradient descent method, where at optimization step $s$ we have:
@@
for some $1>\gamma_{r}>\gamma_{r-1}>\ldots>\gamma_{1}=0$. To understand why this concept is beneficial, let us consider the discretization of the fine-scale problem (i.e., for $\mathcal{G}_{1}$) in Equation (4) using $M$ samples. Suppose that the error in approximating the expected loss at the first, finest scale ${\mathbb{E}}[\mathcal{L}_{s}^{(1)}]$ by the samples behaves as $e^{(1)}={\frac{C}{\sqrt{M}}}$. Now, let us consider a 2-scale process, with $M_{1}$ samples on the finest resolution graph. The error in evaluating the loss on the second scale ${\mathbb{E}}[\mathcal{L}_{s}^{(2)}]$ with $M_{2}$ examples is given by $e^{(2)}={\frac{C}{\sqrt{M_{2}}}}$ while the error in evaluating ${\mathbb{E}}[(\mathcal{L}_{s}^{(1)}-\mathcal{L}_{s}^{(2)})]$ is $e^{(1,2)}={\frac{C\gamma_{1}}{\sqrt{M_{1}}}}$. Since the assumption is that $\gamma<1$, we are able to choose $M_{2}$ and $M_{1}$ such that $e^{(1)}\approx e^{(2)}+e^{(1,2)}$. Hence, rather than using $M$ finest graph resolution observations, we use only $M_{1}<M$ such observations and sample the coarser problems instead.
@@
${\mathbb{E}}\,\mathcal{L}^{(1)}_{s}({\boldsymbol{\theta}})$ $\approx$ $\textstyle{{\frac{1}{M_{R}}}\sum\,\mathcal{L}^{(R)}_{j}({\boldsymbol{\theta}})}+\textstyle{{\frac{1}{M_{R-1}}}\sum\,\left(\mathcal{L}^{(R-1)}_{s}({\boldsymbol{\theta}})-\mathcal{L}^{(R)}_{s}({\boldsymbol{\theta}})\right)}+\textstyle{{\frac{1}{M_{1}}}\sum\,\left(\mathcal{L}^{(1)}_{s}({\boldsymbol{\theta}})-\mathcal{L}^{(2)}_{s}({\boldsymbol{\theta}})\right)},$ (10)
where $M_{R}>M_{R-1}>M_{1}$. This implies that for the computation of the loss on the fine resolution $\mathcal{G}_{1}=\mathcal{G}$, only a few fine graph approximations are needed, and most of the computations are done on coarse, small graphs. Our method is summarized in Algorithm˜2, and we provide an illustration of the computation process in Figure˜3. Naturally, this method suits inductive learning tasks, where multiple graphs are available.
@@
Figure 3: Illustration of the Multiscale Gradient Computation algorithm introduced in Section˜4.3.
@@ 4.4 A Motivating Example
[…]
@@
To explore this further, we train a four-layer GCN (gcn) on a dataset of 100 rods in varying orientations and evaluate different coarsening methods for constructing a reduced graph. Table 1 presents the loss values for various coarsening strategies along with the corresponding edge counts. The results highlight another aspect of our approach: methods that do not increase connectivity prior to pooling (i.e., with $p=1$) may, in certain situations, lead to significant variance in loss across different coarsened graphs. For this dataset, introducing additional connectivity before pooling successfully mitigates this issue while still reducing the graph size in terms of both nodes and edges.
To further evaluate the capabilities of our multiscale methods on this synthetic dataset, Appendix B presents the full multiscale training results and compares them with standard fine-grid training. Additionally, we construct a second synthetic dataset derived from the MNIST image dataset (lecun1998mnist). Results on both datasets demonstrate that our methods effectively capture global graph features while enabling efficient training.
@@ 4.5 Theoretical Properties of Reduced Graphs
At the base of all the above-discussed techniques, the core is the ability of the coarse graph to represent the fine graph. To this end, consider the least squares model problem, in which we aim to minimize over $\boldsymbol{\theta}$} the following loss function:
$\mathcal{L}({\boldsymbol{\theta}})={\frac{1}{2N}}\|\sigma({\bf A}{\bf X}\{{\boldsymbol{\theta}})-{\bf Y}\|^{2},$ (11)
and assume that $\sigma$ is the identity mapping. One way to analyze the solution of the coarse-scale problem is through sketching (see woodruff2014sketching). In sketching, the original problem is replaced by
$\mathcal{L}_{c}({\boldsymbol{\theta}})={\frac{1}{2N}}\|{\bf P}^{\top}{\bf A}{\bf X}{\boldsymbol{\theta}}-{\bf P}^{\top}{\bf Y}\|^{2}$ (12)
where $\bf P$$^{\top}$ is a sketching matrix. We establish the following result with its proof provided in Appendix A, which applies to Sub-to-Full training, and we note is only limited to the linear case $\sigma=Id$:
@@
Let $\boldsymbol{\theta}$$^{*}$ be the solution of equation (11) and let ${\boldsymbol{\theta}}_{C}^{*}$ be the solution of equation (12). Denote $\epsilon>0$, and let $\bf P$$^{\top}$ be a sparse subspace embedding with $O(\frac{c^{2}}{\epsilon})$ rows, where $c$ is the size of the feature vectors (number of graph channels), that injects any vector into a subset of indices, which we denote by $C$. According to this subset, denote ${\bf A}=\begin{bmatrix}{\bf A}_{CC}&{\bf A}_{CF}\\ {\bf A}_{FC}&{\bf A}_{FF}\end{bmatrix}$, and ${\bf X}=\begin{bmatrix}{\bf X}_{C}\\ {\bf X}_{F}\end{bmatrix}$. Then if $\sigma=Id$, there exists $\|{\bf R}\|<\|{\bf A}_{CF}{\bf X}_{F}{\boldsymbol{\theta}}_{C}^{*}\|$ such that with probability .99:
$\mathcal{L}_{c}({\boldsymbol{\theta}}_{C}^{*})={\frac{1}{2N}}\|{\bf P}^{\top}{\bf A}{\bf P}{\bf P}^{\top}{\bf X}{\boldsymbol{\theta}}_{C}^{*}-{\bf P}^{\top}{\bf Y}\|^{2}\leq{\frac{1\pm\epsilon}{2N}}\left(\|{\bf A}{\bf X}{\boldsymbol{\theta}}^{*}-{\bf Y}\|^{2}+\|{\bf R}\|^{2}\right)$ (13)
@@ 5 Computational Complexity
As noted in Section˜3, the computational cost of GNN message-passing layers is directly influenced by the size of both the node feature matrix and the adjacency matrix. This observation suggests that reducing the number of nodes and edges can lead to improved computational efficiency during training.
To illustrate this, consider Graph Convolutional Networks (GCNs) (gcn), where each layer follows the update rule defined in Equation 1. When analyzing floating-point operations (FLOPs), we account for both the input and output dimensions of each layer, as well as the sparsity of the adjacency matrix, meaning that only nonzero entries contribute to computation. The total number of FLOPs required for a single GCN layer is given by:
@@
[…]
Importantly, this effect is not limited to GCNs. A similar computational advantage is observed for other GNN architectures, like Graph Isomorphism Networks (GIN) (gin) and Graph Attention Networks (GAT) (gat). In all cases, reducing the graph size results in fewer computations per layer, demonstrating a generalizable advantage across different GNN architectures.
@@ 6 Experiments
We demonstrate that our approach achieves comparable or superior downstream performance to standard GNN training on the original graph while reducing computational costs, across the training techniques outlined in Section˜4 and the pooling strategies discussed below. We demonstrate our methods on both mid-size and large-scale datasets, showing that given a need to solve a certain graph problem, our methods will prove to be efficient given existing computational resources.
@@
We incorporate various coarsening techniques to enhance GNN training efficiency:
@@
    Topk Coarsening: Selects $n_{r}$ nodes with the highest degrees before computing ${\bf A}_{r}$ using Equation (5). This approach may be computationally expensive for large graphs, as it requires full degree information.
@@
    Subgraph Coarsening: Samples a random node $j$ and selects its nearest $n_{r}$ neighbors. If geometric positions are available, the selection is based on Euclidean distance; otherwise, a khopk-hop ego network is used. This method captures local graph structures but may overlook global properties.
@@
    Hybrid Coarsening: Alternates between Random and Subgraph Coarsening to leverage both local and global components of the graph.
Further details on the coarsening methods are provided in Appendix C.
@@ 6.1 Transductive Learning
We evaluate our approach using the OGBN-Arxiv dataset (ogbn_arxiv) with GCN (gcn), GIN gin, and GAT (gat). This dataset consists of 169,343 nodes and 1,166,243 edges, where nodes represent individual computer science papers, edges correspond to citation relationships where one paper cites another, and each node is labeled with the paper’s primary subject area. We report test accuracy across three runs with different random seeds and include standard deviation to assess robustness, and present our results in Table 2, demonstrating that our multiscale framework achieves comparable or superior performance relative to standard training. Notably, random pooling proves to be the most effective approach for this dataset, offering both efficiency and strong performance. Given the large number of edges in this dataset, we use $p=1$ and report the corresponding edge counts in Table 3. The results highlight the substantial reduction in edges at each pooling level, significantly improving training efficiency. In Table 2, we also provide a timing analysis, including both theoretical FLOP estimates (as elaborated in Section 5) and empirical runtime measurements, confirming that our multiscale approach enhances efficiency while maintaining strong predictive performance. To show the efficiency of our methods, we demonstrate a comparison of the convergence of the training loss between the multiscale training and fine grid training in Appendix D, and we report memory consumption measurements in Appendix E.
@@
We demonstrate our results using the OGBN-MAG dataset (ogbn_arxiv), which is a heterogeneous network composed of a subset of the Microsoft Academic Graph (MAG). This dataset includes papers, authors, institutions, and fields of study, of which the papers are labeled. We focus on the ’paper’ nodes and the ’citing’ relation, using a homogeneous graph dataset of 736,389 nodes and 5,416,271 edges. We summarize our results in Table˜4, where we show that using GCN (gcn), our methods achieve results comparable to the baseline results. We demonstrate our results for 2, 3, and 4 levels of multiscale training, and note that even though GCN is widely used and may be applied to this dataset, other more advanced architectures may yield better results than the presented baseline. However, changing a chosen model will not harm the effectiveness of our proposed methods.
@@
To emphasize that our method is general and will prove to be effective over various datasets and methods, we use additional datasets of various sizes and show that for every one of them, our methods prove to be efficient while yielding comparable or even better results compared to the original baseline. In Appendix F, we show results and details for the well-known Cora, SiteCeer, and PubMed datasets (cora_dataset). We further test our methods on Flickr (flickr), WikiCS (wikics), DBLP (dblp), the transductive versions of Facebook, BlogCatalog, and PPI (facebook_blogcatalog_ppi) datasets, detailed in Appendix G. We present results for Flickr and PPI, along with timing evaluations, as well as full results for all those datasets and additional experimental details. Additionally, in Appendix J, we provide ablation studies, where we study various coarsening ratios and epoch distribution.
@@ 6.2 Inductive Learning
We evaluate our method on the ShapeNet dataset (shapenet), a point cloud dataset with multiple categories, for a node segmentation task. Specifically, we present results on the ’Airplane’ category, which contains 2,690 point clouds, each consisting of 2,048 points. Since this dataset contains multiple graphs, its overall size is large, and achieving an improvement in training efficiency for each batch will yield a significant impact on training efficiency.
In this dataset, nodes are classified into four categories: fuselage, wings, tail, and engine. For training, we use Dynamic Graph Convolutional Neural Network (DGCNN) from dgcnn, which dynamically constructs graphs at each layer using the $k$-nearest neighbors algorithm, with $k$ as a tunable hyperparameter. We test our multiscale approach with random pooling, subgraph pooling, and a combination of both. In subgraph pooling, given that the data resides in $\mathbb{R}^{3}$, we construct subgraphs by selecting a random node and its $n_{j}$ nearest neighbors. Experiments are conducted using three random seeds, and we report the mean Intersection-over-Union (IoU) scores along with standard deviations in the top section of Table 5. Our results show that for all tested values of $k$, applying either random or subgraph pooling improves performance compared to standard training while reducing computational cost. Furthermore, combining both pooling techniques consistently yields the best results across connectivity settings, suggesting that this approach captures both high- and low-frequency components of the graph while maintaining training efficiency. Additional results for multiscale training on the full ShapeNet dataset are provided in Appendix H.
We apply the Multiscale Gradients Computation method to the ShapeNet ’Airplane’ category, evaluating various pooling techniques using Algorithm 2 with two levels. As shown in the bottom section of Table 5, our approach reduces training costs while achieving performance comparable to fine-grid training. These results show that the early stages of optimization can be efficiently performed using an approximate loss, as we propose, with fine-grid calculations needed only once the loss function nears its minimum. Additional results, timing analyses and experimental details for the ShapeNet dataset are provided in Appendix H.
@@
We further evaluate our Multiscale Gradients Computation method on the protein-protein interaction (PPI) dataset (ppi_dataset), consisting of graphs representing different human tissues. The results, with additional details, are provided in Appendix I. Furthermore, Appendix I also presents results on the NCI1 (wale2008comparison) and ogbg-molhiv (ogbn_arxiv) datasets.
7 Limitations
While our methods are broadly applicable across various models, datasets, and architectures, several limitations remain. Although our multiscale approach successfully avoids using the fine graph for most of the training process, a few iterations must still be conducted on the original fine grid, which requires it to fit in memory. Additionally, while the different variants of our method generally yield comparable performance, we currently lack a principled, deterministic criterion for selecting the optimal variant in a given setting, and specific cases may cause relatively high variance in certain datasets and methods. However, in this paper, we focus on the framework and the overall multiscale concept. Lastly, the theoretical analysis presented in this paper is restricted to the linear case (i.e., $\sigma=\text{Id}$). We believe further theoretical investigation may be performed to extend the analysis to nonlinear settings.